40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10680     Entry time: Thu Nov 6 12:53:09 2014
Author: diego 
Type: Update 
Category: ASC 
Subject: X arm restored 

[Diego, Koji]

X arm has been restored, after modifying the two parameters mentioned in http://nodus.ligo.caltech.edu:8080/40m/10676 (C1SUS_ITMX:  LSC/DAMP and LSC/BIAS); after that, a manual re-alignment of ETMX was necessary due to heavy PIT misalignment. I will check the ALS lock once work on the Y arm is done.

ELOG V3.1.3-